# Do LLMs Understand Your Translations? Evaluating Paragraph-level MT with Question Answering

### v2

@@ Abstract
Despite the steady progress in machine translation evaluation, existing automatic metrics struggle to capture how well meaning is preserved beyond sentence boundaries. We posit that reliance on a single intrinsic quality score, trained to mimic human judgments, might be insufficient for evaluating translations of long, complex passages, and a more “pragmatic” approach that assesses how accurately key information is conveyed by a translation in context is needed. We introduce TREQA (Translation Evaluation via Question-Answering), a framework that extrinsically evaluates translation quality by assessing how accurately candidate translations answer reading comprehension questions that target key information in the original source or reference texts. In challenging domains that require long-range understanding, such as literary texts, we show that TREQA is competitive with and, in some cases, outperforms state-of-the-art neural and LLM-based metrics in ranking alternative paragraph-level translations, despite never being explicitly optimized to correlate with human judgments. Furthermore, the generated questions and answers offer interpretability: empirical analysis shows that they effectively target translation errors identified by experts in evaluated datasets. Our code is available at https://github.com/deep-spin/treqa
@@ 1 Introduction
We show that TREQA is competitive with or even outperforms state-of-the-art neural and LLM-based metrics in ranking alternative paragraph-level translations of literary texts translated into English from multiple languages, including Polish, Japanese, Russian, French, German, and Chinese, both in the reference-based and reference-free (or quality-estimation; QE) setups. TREQA performs particularly better when question generation is conditioned on evaluated candidates – a setting that benefits from using more capable LLMs. Furthermore, we demonstrate that the questions generated by TREQA precisely target expert-marked error spans, and generating these specific errors (rather than paraphrasing them) is critical for accurately capturing translation quality differences.
@@ 4 Results
[…]
GembaMQM Sonnet-3.5 51.1 62.0 54.0 61.3
Qwen2.5-72B-IT 55.0 55.3 44.0 55.3
[…]
GPT-4o 65.0 56.7 46.7 43.3 56.0
63.3 56.1 44.7 42.7 52.7
Sonnet-3.5 58.3 55.0 52.0 47.3 58.0
64.4 57.8 44.7 43.3 57.3
TREQA Qwen2.5-72B-IT 64.4 46.7 45.4 46.7 58.7
@@ 6 Related Work
Automatic Evaluation Prior automatic evaluation methods apply sentence-level metrics directly to longer texts or explicitly adapt them for paragraph-level evaluation Deutsch et al. (2023); Raunak et al. (2023); Vernikos et al. (2022), often assuming one-to-one sentence alignment. Such assumptions fail in realistic scenarios involving structural variations (sentence segmentation, reordering, restructuring). Discourse-level metrics Joty et al. (2017); Jiang et al. (2022) address these by evaluating coherence through discourse categories. In contrast, TREQA uses LLM-derived answers to capture deeper semantic and pragmatic nuances beyond span-level alignments.
@@ Acknowledgments
We thank Jessy Li, António Farinhas, Eleftheria Briakou, Duarte Alves and the SARDINE lab team for helpful discussions and feedback on earlier versions of the paper. This work was supported by EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI), and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações.

### v3

@@ Abstract
Despite the steady progress in machine translation evaluation, existing automatic metrics struggle to capture how well meaning is preserved beyond sentence boundaries. We posit that reliance on a single intrinsic quality score, trained to mimic human judgments, might be insufficient for evaluating translations of long, complex passages, and a more “pragmatic” approach that assesses how accurately key information is conveyed by a translation in context is needed. We introduce TREQA (Translation Evaluation via Question-Answering), a framework that extrinsically evaluates translation quality by assessing how accurately candidate translations answer reading comprehension questions that target key information in the original source or reference texts. In challenging domains that require long-range understanding, such as literary texts, we show that TREQA is competitive with and, in some cases, outperforms state-of-the-art neural and LLM-based metrics in ranking alternative paragraph-level translations, despite never being explicitly optimized to correlate with human judgments. Furthermore, the generated questions and answers offer interpretability: empirical analysis shows that they effectively target translation errors identified by experts in evaluated datasets. Our code is available at https://github.com/deep-spin/treqa.
@@ 1 Introduction
We evaluate TREQA on literary texts – an especially challenging and underexplored domain for MT evaluation as these texts demand deep semantic understanding, multi-sentence reasoning, and subtle discourse-enforced meaning nuances that traditional metrics would fail to capture. We show that TREQA is competitive with or even outperforms state-of-the-art neural and LLM-based metrics in ranking alternative paragraph-level translations of literary texts translated into English from multiple languages, including Polish, Japanese, Russian, French, German, and Chinese, both in the reference-based and reference-free (or quality-estimation; QE) setups. TREQA performs particularly better when question generation is conditioned on evaluated candidates – a setting that benefits from using more capable LLMs. While literature is open to interpretation, many questions generated by TREQA focus on elements such as character relationships, event sequences, and causal links, which have stable, answerable properties across translations. Furthermore, we demonstrate that the questions generated by TREQA precisely target expert-marked error spans, and generating these specific errors (rather than paraphrasing them) is critical for accurately capturing translation quality differences.
@@ 3 Experimental Setup
Significance Testing We report clusters for metrics based on statistically significant performance gaps. Specifically, we conduct pairwise Wilcoxon signed-rank tests between all metric pairs within each group, where the dataset and the type of evaluation (Ref or QE) define groups. To account for multiple comparisons, we apply the Benjamini-Hochberg FDR correction to the resulting p-values. Based on the corrected p-values (thresholded at 0.05), we construct a significance matrix indicating which metric differences are statistically significant. This matrix serves two purposes: identifying metrics that are significantly better than others for ranking, and acting as a distance matrix for agglomerative clustering to group metrics with similar significance patterns, resulting in final ranking.
@@ 4 Results
[0.5pt/2pt] Comet 62.2 1
[0.5pt/2pt] CometQE 46.1 4 37.3 3 33.3 2 51.3 1
[0.5pt/2pt] xComet 61.7 2 55.6 3 53.3 2 41.3 2 54.7 1
[0.5pt/2pt] MTEQA 43.3 3
[0.5pt/2pt] GPT-4o 40.6 4 44.0 3 33.3 2 52.0 1
GembaMQM Sonnet-3.5 51.1 4 62.0 1 54.0 1 61.3 1
Qwen2.5-72B-IT 55.0 3 55.3 2 44.0 2 55.3 1
[0.5pt/2pt] 59.4 2 43.9 4 44.0 3 40.0 2 54.0 1
GPT-4o 65.0 1 56.7 2 46.7 3 43.3 2 56.0 1
63.3 1 56.1 3 44.7 3 42.7 2 52.7 1
Sonnet-3.5 58.3 2 55.0 3 52.0 2 47.3 2 58.0 1
64.4 1 57.8 1 44.7 3 40.0 2 57.3 1
TREQA Qwen2.5-72B-IT 64.4 1 46.7 4 45.4 3 46.7 2 58.7 1
@@ 6 Related Work
Automatic Evaluation Prior automatic evaluation methods apply sentence-level metrics directly to longer texts or explicitly adapt them for paragraph-level evaluation (Deutsch et al., 2023; Raunak et al., 2023; Vernikos et al., 2022), often assuming one-to-one sentence alignment. Such assumptions fail in realistic scenarios involving structural variations (sentence segmentation, reordering, restructuring). Discourse-level metrics (Joty et al., 2017; Jiang et al., 2022) address these by evaluating coherence through discourse categories. In contrast, TREQA uses LLM-derived answers to capture deeper semantic and pragmatic nuances beyond span-level alignments.
@@
7 Practical Recommendations
Our experiments offer several practical insights for those looking to use TREQA to evaluate paragraph-level MT, particularly in complex domains like literary texts as discussed below:
  • Start with Candidate-Aware QAG: Conditioning the generation of questions on candidate translations to be evaluated along with the source / reference texts (the “with candidates” setting) generally yields the best results, as it helps the LLM focus on discriminative differences and potential errors (Sections 4, 5, D.1).
  • Use Powerful LLMs: Leveraging the most capable LLMs available for both Question-Answer Generation and Question Answering is crucial. Although there is no single best model across all scenarios, models like GPT-4o, Sonnet-3.5, and Qwen2.5-72B-IT consistently perform well (Table 1, Figure 3).
  • Always Regenerate Answers: Re-generating answers using the QA model, grounded in the source/reference text, is essential to mitigate strong positional biases and hallucinations and ensure that the evaluation truly reflects comprehension of the ground truth rather than copying from candidates (Figure 5).
  • Consider Sampling for QAG: While greedy decoding offers consistency, using sampling (e.g., temperature > 0) during question generation can explore a wider range of potential issues and often leads to higher mean accuracy over multiple runs, especially with larger models (Figure 3).
  • Non-English Target Texts: The current TREQA framework relies on English questions and shows reduced QA performance when the context is not in English (D.4). Hence for evaluating translations into languages other than English, we would recommend leveraging newer multilingual LLMs with stronger cross-lingual QA capabilities with TREQA.
  • Iterate and Inspect: The questions generated by TREQA offer interpretability. Manually inspecting the questions and answers can provide insights into why a translation is favored and can help refine prompts or settings for specific use cases.
@@ Acknowledgments
We thank Jessy Li, António Farinhas, Eleftheria Briakou, Duarte Alves, and the SARDINE lab team for helpful discussions and feedback on earlier versions of the paper. This work was supported by EU’s Horizon Europe Research and Innovation Actions (UTTER, contract 101070631), by the project DECOLLAGE (ERC-2022-CoG 101088763), by the Portuguese Recovery and Resilience Plan through project C645008882-00000055 (Center for Responsible AI), and by FCT/MECI through national funds and when applicable co-funded EU funds under UID/50008: Instituto de Telecomunicações.